# On Efficient Variants of Segment Anything Model: A Survey

### v1

@@ 2.3.4 Low-Rank Factorization
Low-rank factorization is the technique that decomposes the high dimension matrix into the product of low dimension matrices, aiming at reducing parameters of models for less space occupation and higher computational efficiency. A commonly used method of low-rank factorization is the singular value decomposition (SVD). It factorizes the complex matrix $A\in\mathbb{R}^{m\times n}$ into three matrices and the calculation can be further rewritten as follows:
@@ 3 Efficient Variants of SAM
This section reviews the efforts to develop lightweight, efficient SAM-like models that have emerged since SAM gained prominence. These works aim to reduce the model’s high computational cost and enable efficient performance, while preserving SAM’s robust segmentation capabilities and generalization strength. As outlined in Section 2.1.2, SAM addresses two primary tasks including Segment Anything (SegAny) and Segment Everything (SegEvery). Accordingly, we discuss the research aimed at improving each task separately: Section 3.1 focuses on accelerating the SegAny task, and Section 3.2 covers efforts to accelerate the SegEvery task. Notably, some methods are applicable to both tasks, and we discuss these contributions individually. Additionally, we categorize all models into different classes based on the techniques they employ, and the taxonomy of methodologies is presented in Fig. 3. Finally, in Section 3.3, we outline four potential directions for future research in this area.
@@ 3.1.1 Training from Scratch
Figure 4: The architecture of FastSAM [161]. It takes two stages to achieve segmenting anything: all-instance segmentation (the upper part) and prompt-guided selection (the lower part). It is worth noting that FastSAM have introduced CLIP[89] to support text prompt.
@@ 3.1.2 Knowledge Distillation based Methods
Similar to MobileSAM, the later proposed ESAM [162] utilizes EfficientFormerV2 [62] as its backbone, aiming for improved performance in CPU environments, particularly resource-constrained medical devices. Given that expert models [125, 120, 12] often outperform SAM in medical applications, ESAM incorporates a novel Knowledge Distillation (KD) strategy called Holistic Knowledge Distillation (HKD) to transfer knowledge from an expert model to ESAM. HKD involves two components: distillation on feature maps and distillation on output masks. For the feature map distillation, three different methods with varying focuses [13, 102, 138] are combined to guide the learning process. For the output mask distillation, ESAM uses the Mean Square Error (MSE) loss between the teacher’s and student’s masks, supplemented by a Binary Cross-Entropy (BCE) loss between the teacher’s masks and the ground truth masks. To further align the feature maps between the expert model and ESAM, a Teacher Guided Module (TGM) is introduced as illustrated in Fig. 9.
Figure 9: The framework of distilling ESAM [162] from expert model. ESAM adopts the EfficientFormerV2 [62] as image encoder and a Teacher Guided Module (TGM) is applied to align the feature maps for holistic konwledge distillation (HKD).
@@
Shen et al. [101] introduced FastSAM3D, an efficient segment-anything model specifically designed for 3D volumetric medical images. The key contribution of this work is the development of the 3D Sparse Flash Attention mechanism. This novel attention approach combines the advantages of the 3D Dilated Attention [26], which extends the receptive field, with FlashAttention [23] to accelerate computations. FastSAM3D utilizes a modified ViT-Tiny as the image encoder, distilled from the ViT-Base encoder, ensuring efficiency without compromising performance. The authors implemented a layer-wise progressive distillation strategy to iteratively align feature maps between the two encoders.
@@
Figure 12: The overview of RWKV-SAM [145] (left) and its image encoder (right). The image encoder is made up of three stages, where the first two stages use Mobile Convolution Blocks [97] and the last stage uses Vision RWKV blocks [30].
A recent work by Yuan et al. [145], RWKV-SAM, represents a significant advancement in accelerating SAM by introducing the popular linear attention model [39, 87] as an efficient backbone. In their study, they compare RWKV-based and Mamba-based architectures and select the RWKV-based approach to construct a lightweight version of SAM. The backbone is a hybrid design, with the first two stages consisting of Mobile Convolution Blocks from [97], and the final stage built using Vision RWKV blocks [30]. More details on RWKV can be found in Section 2.2.2. Additionally, a refinement module is incorporated into the SAM-like architecture to enhance mask quality by fusing features from different levels generated at each stage. The overall architecture of RWKV-SAM is depicted in Fig. 12. The model is trained using a ”distillation-finetune” strategy, where knowledge from SAM-H is first distilled into the backbone, followed by fine-tuning of the entire model. RWKV-SAM demonstrates significant improvements in efficiency, while maintaining comparable segmentation performance to SAM.
@@ 3.2 Accelerating SegEvery Tasks
Figure 14: Comparison between the default sampling strategy (a-b) for SAM and the hierarchical sampling strategy (c-e) for TinySAM [103]. The hierarchical strategy contains two rounds of sampling: one round for sparse sampling (c) and the other round for dense sampling (d). The final prediction comes from the fusion of results in two rounds.
@@ 3.3 Future Research Directions
Efficient Segmentation for Video and Multi-Modality Data. Video and multi-modality tasks, which deal with complex, dynamic environments, are rapidly gaining relevance across numerous real-world applications. While some initial efforts, such as SAM 2 [93] for video segmentation and MM-SAM [128] for multi-modality tasks, have extended SAM’s applicability, efficiency remains a pressing issue. Video data contains temporal redundancy, while multimodal data often exhibits correlations between modalities. Leveraging these inherent redundancies through techniques like temporal aggregation and cross-modal feature sharing could significantly reduce computational costs. Future work could focus on optimizing SAM’s runtime complexity by utilizing spatiotemporal attention, efficient memory mechanisms for temporal data, and early-fusion techniques to reduce the number of modality-specific computations. Developing frameworks that adapt dynamically to different levels of redundancy across frames or modalities can further drive computational efficiency in real-world applications.
@@ 4 Evaluation
In this section, we systematically compare the efficiency and accuracy of the previously described SAM variants. With reference to the experiments conducted in these works, we choose the tasks that most of the works have carried out and evaluate them on their commonly used datasets with corresponding metrics. Our evaluations are conducted on a single 24GB RTX 3090 GPU with a 14 vCPU Intel(R) Xeon(R) Gold 6330 Processor @ 2.00GHz. More details are provided in the following subsections: Section 4.1 introduces the datasets and metrics used for evaluation; Sections 4.2 and 4.3 report the quantitative results of efficiency and accuracy, respectively.
@@ 4.1 Datasets and Metrics
We select COCO 2017 [65] and LVIS v1 [40] as the evaluation datasets. COCO is a large-scale dataset designed for object detection, segmentation, and captioning, containing 330K images and 1.5 million object instances across 80 object categories. LVIS is tailored for large vocabulary instance segmentation, featuring over 2 million high-quality segmentation masks across more than 1200 categories in 164K images. For our evaluation, we use the validation sets of both datasets, which include 36,781 instances in 5,000 images for COCO and 244,707 instances in 19,809 images for LVIS. To evaluate efficiency, we first test several soft indicators like #$\#Params$, $FLOPs$, $MACs$, and $Memory\ Usage$. And we further calculate the Efficient Error Rate ($EER$) a more comprehensive evaluation, as outlined in [86]. EER is defined as,
@@
where $N$ is the number of metrics and $M_{i}$, $R_{i}$ refer to the i-th metric of the tested model and the reference one, respectively. In our evaluation, we set the reference model to be SAM-H. In addition to these metrics, we also report the runtime and throughput of the models. For accuracy evaluation, we use mean intersection over union ($mIoU$) for the SegAny task and mean average precision ($AP$) for instance segmentation.
@@ 4.2 Efficiency Comparison
We first report the efficiency results of SAM and its variants. With the picture groceries.jpg [^1] https://github.com/facebookresearch/segment-anything/blob/main/notebooks/images/groceries.jpg used in SAM’s official example as input, we utilize a bounding box as prompt and evaluate models’ #$\#Params$, $FLOPs$ and $MACs$ by leveraging the tool calflops [^2] https://github.com/MrYxJ/calculate-flops.pytorch. We also calculate the $ERR$ for comprehensive comparison. The results are illustrated in Tab. 2. Among the efficient variants, we observe that EdgeSAM has the lowest number of parameters, $FLOPs$, $MACs$ and consequently the $EER$, while EfficientViT-SAM-XL1 takes the highest numbers whose $EER$ is 3% more than SAM-B. When compared to the heaviest SAM-H, all variants present obvious reduction in both model size and calculation.
Model #Params GFLOPs GMACs EER (%)
SAM-H 641.09M 5490 2730 100
SAM-L 312.34M 2640 1310 48.26
SAM-B 93.74M 746.4 370.47 13.93
SAM2-B+ 80.83M 533.87 266.44 10.70
FastSAM 68M 887.6 - 13.39
MobileSAM 10.13M 76.39 37.49 1.45
EdgeSAM 9.58M 44.02 21.81 1.03
EfficientSAM-Ti 10.22M 53.64 26.74 1.18
EfficientSAM-S 26.41M 371.68 185.56 5.90
RepVit-SAM 27.22M 231.65 115.07 4.23
EfficientViT-SAM-L0 34.79M 154.24 76.89 3.68
EfficientViT-SAM-XL1 203.35M 523.07 261.21 16.94
SlimSAM-50 28.01M 196.21 96.73 3.83
SlimSAM-77 9.85M 47.66 23.19 1.08
TinySAM 10.13M 76.39 37.94 1.45
Table 2: Quantitative results of #Params, GFLOPs, GMACs and EER. P.S. The results of FastSAM [161] are collected from [164] and others are evaluated under our framework.
@@
We also measure the models’ inference time in both SegAny and SegEvery modes, using 100 images from the COCO validation set as evaluation data. For the SegAny task, each image is prompted with 50 fixed bounding boxes. We report the cumulative time for every 10 boxes through a curve chart (shown in Fig. 16). Based on this, we calculate the average time required to process one image with one box prompt and report it as the inference time for the SegAny task. The evaluation is conducted on both CPU and GPU environments, and we simultaneously record GPU memory usage. Additionally, we test the throughput of each variant on the COCO validation set, using the ground truth boxes. The results are summarized in Tab. 3. Our findings reveal that EfficientViT-SAM-L0 achieves the shortest inference time for the SegAny task, offering nearly 30x acceleration on GPU and almost 50x acceleration on CPU compared to the heaviest model, SAM-H. EdgeSAM also shows impressive performance with a low CPU latency of 259 ms, while NanoSAM achieves a 20 ms latency on GPU, both coming close to the best results. In the throughput test on the COCO dataset, NanoSAM leads with 27.9 images processed per second. Two other variants, EfficientSAM-Ti and EfficientViT-SAM-L0, also demonstrate strong throughput, each exceeding 20 images per second.
Model Memory (GB) Latency-CPU (ms) Latency-GPU (ms) Throughput (img/s)
SAM-H 7.46 9470 472 2.04
SAM-L 6.03 6800 289 3.33
SAM-B 4.39 3294 129 6.78
SAMfast-H 2.13 - 127 6.36
SAM2-B+ 1.38 1221 46 13.82
FastSAM 4.96 850 103 8.42
MobileSAM 1.81 424 44 16.79
EdgeSAM 1.72 259 49 17.56
EfficientSAM-Ti 4.46 1159 61 20.04
EfficientSAM-S 7.53 2605 86 19.7
RepVit-SAM 1.87 1013 59 13.49
EfficientViT-SAM-L0 1.73 194 16 20.92
EfficientViT-SAM-XL1 2.57 1334 67 12.28
SlimSAM-50 4.49 2312 87 9.13
SlimSAM-77 4.34 2230 73 10.28
TinySAM 2.2 422 42 16.53
Q-TinySAM 2.34 697 73 7.03
NanoSAM 0.94 - 20 27.9
Table 3: Quantitative results of inference latency of SegAny task and throughput on COCO dataset. P.S. 1) We also report the GPU memory usage in first column when models are inferring on GPU; 2) SAMfast [88] and NanoSAM [84] can not run on CPU environment.
Model 16*16 grid (ms) 32*32 grid (ms) 64*64 grid (ms) other prompt strategies (ms)
@@
Table 4: Quantitative results of inference latency of SegEvery task with grid sampling in different scales or other efficient strategies. P.S. OOM denotes out-of-memory.
For the SegEvery task, we report the mean time required to generate all masks for an image using different point grid sizes (16*16, 32*32, 64*64) or specialized sampling strategies. The results are presented in Tab. 4. With the default 32*32 grid, SAMfast-H demonstrates the highest efficiency, with a latency of 848 ms—more than twice as fast as SAM-H. EfficientViT-SAM-L0 performs best with the 16*16 and 64*64 grids, achieving latencies of 258 ms and 3938 ms, respectively. Interestingly, we observe that EfficientSAM-S is slower than SAM-H when using lower grid densities, showing latencies of 1100 ms for the 16x16 grid and 2290 ms for the 32*32 grid. Models employing alternative sampling strategies demonstrate significant improvements in efficiency, particularly FastSAM, which records a latency of 196 ms, and MobileSAMv2, with a latency of 173 ms.

### v2

@@ 2.3.4 Low-Rank Factorization
Low-rank factorization is the technique that decomposes the high-dimension matrix into the product of low-dimension matrices, aiming at reducing the parameters of models for less space occupation and higher computational efficiency. A commonly used method of low-rank factorization is the singular value decomposition (SVD). It factorizes the complex matrix $A\in\mathbb{R}^{m\times n}$ into three matrices and the calculation can be further rewritten as follows:
@@ 3 Efficient Variants of SAM
This section reviews the efforts to develop lightweight, efficient SAM-like models that have emerged since SAM gained prominence. These works aim to reduce the model’s high computational cost and enable efficient performance, while preserving SAM’s robust segmentation capabilities and generalization strength. As outlined in Section 2.1.2, SAM addresses two primary tasks including Segment Anything (SegAny) and Segment Everything (SegEvery). Accordingly, we discuss the research aimed at improving each task separately: Section 3.1 focuses on accelerating the SegAny task, and Section 3.2 covers efforts to accelerate the SegEvery task. Notably, some methods are applicable to both tasks, and we discuss these contributions individually. Additionally, we categorize all models into different classes based on the techniques they employ, and the taxonomy of methodologies is presented in Fig. 3. Finally, in Section 3.3, we outline the potential directions for future research in this area.
@@ 3.1.1 Training from Scratch
Figure 4: The architecture of FastSAM [173]. It takes two stages to achieve segmenting anything: all-instance segmentation and prompt-guided selection. It is worth noting that the outputs of first stage are used directly as the SegEvery results.
@@ 3.1.2 Knowledge Distillation based Methods
Similar to MobileSAM, the later proposed ESAM [174] utilizes EfficientFormerV2 [69] as its backbone, aiming for improved performance in CPU environments, particularly resource-constrained medical devices. Given that expert models [137, 132, 12] often outperform SAM in medical applications, ESAM incorporates a novel Knowledge Distillation (KD) strategy called Holistic Knowledge Distillation (HKD) to transfer knowledge from an expert model to ESAM. HKD involves two components: distillation on feature maps and distillation on output masks. For the feature map distillation, three different methods with varying focuses [13, 114, 150] are combined to guide the learning process. For the output mask distillation, ESAM uses the Mean Square Error (MSE) loss between the teacher’s and student’s masks, supplemented by a Binary Cross-Entropy (BCE) loss between the teacher’s masks and the ground truth masks. A Teacher Guided Module (TGM) is introduced, to further align the feature maps between the expert model and ESAM.
@@
Shen et al. [113] introduced FastSAM3D, an efficient segment-anything model specifically designed for 3D volumetric medical images. The key contribution of this work is the development of the 3D Sparse Flash Attention mechanism. This novel attention approach combines the advantages of the 3D Dilated Attention [28], which extends the receptive field, with FlashAttention [25] to accelerate computations. FastSAM3D utilizes a modified ViT-Tiny as the image encoder, distilled from the ViT-Base encoder, ensuring efficiency without compromising performance. The authors implemented a layer-wise progressive distillation strategy to iteratively align feature maps between the two encoders, as shown in Fig. 11.
Figure 11: The framework of layer-wise progressive distillation of FastSAM3D [113]. It takes the ViT-B as teacher model to distill the efficient image encoder, a 6-layer ViT-Tiny variant.
@@
A recent work by Yuan et al. [157], RWKV-SAM, represents a significant advancement in accelerating SAM by introducing the popular linear attention model [43, 99] as an efficient backbone. In their study, they compare RWKV-based and Mamba-based architectures and select the RWKV-based approach to construct a lightweight version of SAM. The backbone is a hybrid design, with the first two stages consisting of Mobile Convolution Blocks from [109], and the final stage built using Vision RWKV blocks [32]. More details on RWKV can be found in Section 2.2.2. Additionally, a refinement module is incorporated into the SAM-like architecture to enhance mask quality by fusing features from different levels generated at each stage. The model is trained using a ”distillation-finetune” strategy, where knowledge from SAM-H is first distilled into the backbone, followed by fine-tuning of the entire model. RWKV-SAM demonstrates significant improvements in efficiency, while maintaining comparable segmentation performance to SAM.
@@ 3.2 Accelerating SegEvery Tasks
Figure 13: The pipeline of hierarchical sampling strategy for TinySAM [115]. The hierarchical strategy contains two rounds of sampling: one round for sparse sampling and the other round for dense sampling (including removing points from high confident regions). The final prediction comes from the fusion of results in two rounds.
@@ 3.3 Future Research Directions
Computation Cost Attacks and Defenses. Substantial progress has been made by the research community to improve the efficiency of deep learning models, particularly for deployment on resource-constrained edge devices. However, recent studies have exposed vulnerabilities in some models when subjected to computation cost attacks, which deliberately increase resource consumption and processing time far beyond normal operational levels [52, 14, 15, 38, 97]. Despite these findings, the efficiency and robustness of Segment Anything Model (SAM) variants under computation cost attacks remain largely unexplored, especially when techniques like dynamic models are applied. This presents a valuable opportunity for future research. One direction could involve identifying specific vulnerabilities in SAM variants when exposed to such attacks. Additionally, the insights gained from these vulnerabilities could inform the development of defense strategies that ensure stable efficiency, particularly in resource-limited environments such as edge and mobile devices.
Efficient Segmentation for Video and Multi-Modality Data. Video and multi-modality tasks, which deal with complex, dynamic environments, are rapidly gaining relevance across numerous real-world applications. While some initial efforts, such as SAM 2 [105] for video segmentation and MM-SAM [140] for multi-modality tasks, have extended SAM’s applicability, efficiency remains a pressing issue. Video data contains temporal redundancy, while multimodal data often exhibits correlations between modalities. Leveraging these inherent redundancies through techniques like temporal aggregation and cross-modal feature sharing could significantly reduce computational costs. A recent work, named Surgical SAM 2 [77], has proposed the efficient frame pruning strategy which only retains the most informative frames in memory bank to achieve reduction in memory usage and acceleration in inference, providing a promising idea for efficient segmentation in videos. Future work could focus on optimizing SAM’s runtime complexity by utilizing spatiotemporal attention, efficient memory mechanisms for temporal data, and early-fusion techniques to reduce the number of modality-specific computations. Developing frameworks that adapt dynamically to different levels of redundancy across frames or modalities can further drive computational efficiency in real-world applications.
@@ 4 Evaluation
In this section, we systematically compare the efficiency and accuracy of the previously described SAM variants. With reference to the experiments conducted in those works, we choose the tasks that most of the works have carried out and evaluate them on their commonly used datasets with corresponding metrics. Unless otherwise stated, our evaluations are conducted on a single 24GB RTX 3090 GPU with a 14 vCPU Intel(R) Xeon(R) Gold 6330 Processor @ 2.00GHz. For fair comparisons, we ensure that all models are tested under the same environments and that all tested images are retained at their original resolutions. More details are provided in the following subsections: Section 4.1 introduces the datasets and metrics used for evaluation; Sections 4.2 and 4.3 report the quantitative results of efficiency and accuracy, respectively.
@@ 4.1 Datasets and Metrics
We utilize COCO 2017 [74] and LVIS v1 [45] as the evaluation datasets. COCO is a large-scale dataset for object detection, segmentation, and captioning, containing 330K images and 1.5 million object instances across 80 object categories. LVIS is tailored for large vocabulary instance segmentation, featuring over 2 million high-quality segmentation masks across more than 1200 categories in 164K images. For our evaluation, we use the validation sets of both datasets, which include 36,781 instances in 5,000 images for COCO and 244,707 instances in 19,809 images for LVIS. To evaluate efficiency, we first test several soft indicators like #$\#Params$, $FLOPs$, and $Memory\ Usage$. And we further calculate the Efficient Error Rate ($EER$) for a more comprehensive evaluation, as outlined in [98]. $EER$ is defined as,
@@
where $N$ is the number of metrics and $M_{i}$, $R_{i}$ refer to the i-th metric of the tested model and the reference one, respectively. In our evaluation, we set the reference model to be SAM-H. In addition to these metrics, we also report the runtime and throughput of these models. For accuracy evaluation, we use mean intersection over union ($mIoU$) for the SegAny task and mean average precision ($AP$) for instance segmentation.
@@ 4.2 Efficiency Comparison
We first report the efficiency results of SAM and its variants. With the picture groceries.jpg [^1] https://github.com/facebookresearch/segment-anything/blob/main/notebooks/images/groceries.jpg used in SAM’s official example as input, we utilize a bounding box as prompt and evaluate models’ #$\#Params$ and $FLOPs$ by leveraging the tool calflops [^2] https://github.com/MrYxJ/calculate-flops.pytorch. We then calculate the $EER$ for further comparison. The results are illustrated in Tab. 2. Among the efficient variants, we observe that EdgeSAM has the lowest number of parameters, $FLOPs$ and consequently the $EER$, while EfficientViT-SAM-XL1 takes the highest numbers whose $EER$ is even 6.51% more than SAM-B. When compared to the heaviest SAM-H, all variants present obvious reduction in both model size and calculation.
[…]
Model Params FLOPs EER
SAM-H 641.09 5490 100%
SAM-L 312.34 2640 48.40%
SAM-B 93.74 746.4 14.11%
SAM2-B+ 80.83 533.87 11.17%
FastSAM* 68 887.6 13.39%
MobileSAM 10.13 76.39 1.49%
EdgeSAM 9.58 44.02 1.15%
EfficientSAM-Ti 10.22 53.64 2.57%
EfficientSAM-S 26.41 371.68 5.44%
RepVit-SAM 27.22 231.65 4.23%
EfficientViT-SAM-L0 34.79 154.24 4.12%
EfficientViT-SAM-XL1 203.35 523.07 20.62%
SlimSAM-50 28.01 196.21 3.97%
SlimSAM-77 9.85 47.66 1.20%
TinySAM 10.13 76.39 1.49%
@@
Table 3: Quantitative results of inference latency (ms) and of SegAny task, throughput (img/s) on COCO validation dataset, and the approximation of GPU memory usage (GB). More details are described in Section 4.2. For the latency comparison, we report the results on a CPU-only environment, various GPU-based environments (1080Ti/11GB, 2080Ti/11GB, 3090/24GB, and A800/80GB), and the edge device (Jetson Nano 4GB). The throughput is measured with a single 3090 GPU.
Latency
Model Memory CPU Nano 1080Ti 2080Ti 3090 A800 Throughput
SAM-H 7.46 9470 - 1039 711 461 431 2.04
SAM-L 6.03 6800 - 603 400 270 248 3.33
SAM-B 4.39 3294 - 270 170 116 103 6.78
SAMfast-H 2.13 - - - 939 127 57 6.36
SAM2-B+ 1.38 1221 - 197 110 85 61 13.82
FastSAM 4.96 850 - 183 142 103 53 8.42
MobileSAM 1.81 424 1328 60 38 30 25 16.79
EdgeSAM 1.72 259 713 40 32 24 24 17.56
EfficientSAM-Ti 4.46 1159 - 92 51 40 30 20.04
EfficientSAM-S 7.53 2605 - 188 100 73 60 19.7
RepVit-SAM 1.87 1013 2685 104 69 44 38 13.49
EfficientViT-SAM-L0 1.73 194 651 30 23 16 15 20.92
EfficientViT-SAM-XL1 2.57 1334 3779 116 82 52 36 12.28
SlimSAM-50 4.49 2312 - 192 116 76 68 9.13
SlimSAM-77 4.34 2230 - 161 94 65 55 10.28
TinySAM 2.2 422 1322 60 39 29 25 16.53
Q-TinySAM 2.34 697 2217 116 78 58 48 7.03
NanoSAM 0.94 - 345 39 26 20 16 27.9
Table 4: Quantitative results of inference latency (ms) of SegEvery task with grid sampling in different scales or other efficient strategies. More details are reviewed in Section 3.2.
Grid Sampling Other
Model $16^{2}$ $32^{2}$ $64^{2}$
@@
[…]
For the SegEvery task, we report the mean time required to generate all masks for an image using different point grid sizes (16*16, 32*32, 64*64) or specialized sampling strategies (detailed in Section 3.2). The results are presented in Tab. 4. With the default 32*32 grid, SAMfast-H demonstrates the highest efficiency, with a latency of 848 ms—more than twice as fast as SAM-H. EfficientViT-SAM-L0 performs best with the 16*16 and 64*64 grids, achieving latencies of 258 ms and 3938 ms, respectively. Interestingly, we observe that EfficientSAM-S is slower than SAM-H when using lower grid densities, showing latencies of 1100 ms for the 16*16 grid and 2290 ms for the 32*32 grid. Models employing alternative sampling strategies demonstrate significant improvements in efficiency, particularly FastSAM, which records a latency of 106 ms, and MobileSAMv2, with a latency of 173 ms.